# Notions of entropy for unitary representations

### v2

@@ 4.1 Tuples of vectors and completely positive maps
Let $\pi$ be a representation of $\mathfrak{A}$. Let $v_{1}$, …, $v_{k}\in H_{\pi}$, and regard the tuple $V:=[v_{1},\dots,v_{k}]$ as a linear map from $\mathbf{C}^{\oplus k}$ to $H_{\pi}$. To keep track of how these vectors move relative to each other under the action of $\pi$, we can consider the $\mathrm{M}_{k}$-valued function
@@
If $\varphi$ is a continuous linear map from $\mathfrak{A}$ to $\mathrm{M}_{k}$, then a classic theorem of Stinespring provides necessary and sufficient conditions for it to equal $\Phi^{\pi}_{V}$ for some representation $\pi$ and $k$-tuple $V$: this holds if and only if $\varphi$ is completely positive. See, for example, [Sti55] or [Con00, Theorem 34.7]. The most substantial part of this result is the construction of $\pi$ and $V$ from $\varphi$ if it is completely positive. This is an $\mathrm{M}_{k}$-valued generalization of the GNS construction. If we require in addition that $V$ be cyclic in $\pi$, then the resulting pair $(\pi,V)$ is unique up to unitary equivalence. It is called the minimal dilation of $\varphi$, and we denote its underlying representation by $\pi_{\varphi}$.
In the present paper we do not use the property of complete positivity directly, but we adopt this as the standard term for the class of maps we need to work with. We write $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ for the space of all continuous linear maps from $\mathfrak{A}$ to $\mathrm{M}_{k}$, and $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ for the subset of all completely positive (not just positive) ones.
Generalizing our usage in the scalar-valued case, we write that the completely positive map in (4.1) is associated to $\pi$ by $v_{1}$, …, $v_{k}$, or we call $\Phi^{\pi}_{V}$ the type of the tuple $V$ in $\pi$.
A completely positive map $\varphi:\mathfrak{A}\to\mathrm{M}_{k}$ is called normalized if $\overline{\mathrm{tr}}_{k}\varphi(1)=1$, and unital if $\varphi(1)=I_{k}$. A tuple of vectors in $H_{\pi}$ is orthonormal if and only if their type is unital.
@@
As a vector space topology, the weak topology on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ also defines a canonical uniform structure on this space (see, for instance, [Eng89, Section 8.1] for the basics of uniform structures). Restricted to any bounded subset of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$, this uniform structure is generated by any translation-invariant metric compatible with the topology. We need this uniform structure a few times through the following lemma:
@@
These properties are elementary for the inner product map $H_{\pi}\times H_{\pi}\to\mathbf{C}$. They follow for types by arguing pointwise for each $i$, $j$, and $a\in\mathfrak{A}$. ∎
The results about positive functionals recalled in Section 3 have standard generalizations to $\mathrm{M}_{k}$-valued completely positive maps, and the proofs generally run along the same lines. Alternatively, many facts about $\mathrm{M}_{k}$-valued completely positive maps can be reduced directly to the scalar-valued case by regarding tuples of vectors in a representation from an alternative point of view. We explain this in the next two subsections.
@@ 4.2 Matrices over algebras
If $\mathfrak{A}$ is a C-algebra and $k$ is a positive integer, then we write $\mathrm{M}_{k}(\mathfrak{A})$ for the algebra of $k$-by-$k$ matrices with entries from $\mathfrak{A}$. The algebra operations combine those of $\mathfrak{A}$ with the usual rules for matrices, and we define an involution on $\mathrm{M}_{k}(\mathfrak{A})$ by transposing and applying the involution of $\mathfrak{A}$ entry-wise. Then $\mathrm{M}_{k}(\mathfrak{A})$ it is a separable C*-algebra in a canonical way. It has a natural identification with $\mathfrak{A}\otimes\mathrm{M}_{k}$, and we pass freely between these two points of view; see, for instance, [Con00, Section 34].
If $\pi$ is a representation of $\mathfrak{A}$ on $H$ and $k$ is a positive integer, then we define a representation $\pi^{(k)}$ of $\mathrm{M}_{k}(\mathfrak{A})$ on $H^{\oplus k}$ by combining $\pi$ with the usual rules of matrix-vector multiplication:
$\pi^{(k)}([a_{ij}])\cdot\left[\begin{array}[]{c}\vdots\\ v_{i}\\ \vdots\end{array}\right]:=\left[\begin{array}[]{c}\vdots\\ \sum_{j=1}^{k}\pi(a_{ij})v_{j}\\ \vdots\end{array}\right].$
From another point of view, we can identify $H^{\oplus k}$ with $H\otimes\mathbf{C}^{\oplus k}$, and then $\pi^{(k)}$ is the Kronecker product of $\pi$ with the canonical representation
@@
    We have $\pi^{(k)}\otimes I_{\ell}\cong(\pi\otimes I_{\ell})^{(k)}$ for every positive integer $k$ and {${}\ell\in\{1,2,\dots,\}\cup\{\infty\}$.
@@
Any representation $\kappa$ of $\mathrm{M}_{k}(\mathfrak{A})\cong\mathfrak{A}\otimes\mathrm{M}_{k}$ on a Hilbert space $K$ is generated by commuting representations $\kappa|\mathrm{M}_{k}$ and $\kappa|\mathfrak{A}$. For $\mathrm{M}_{k}$, every representation is an inflation of $\mathrm{canon}_{k}$: this is a classical result of pure algebra, or a special case of the representation theory of C*-algebras of compact operators [Con00, Section 16]. We can therefore write $K$ as $H^{(k)}$ for some auxiliary Hilbert space $H$ so that $\kappa|\mathrm{M}_{k}$ is identified with $I_{H}\otimes\mathrm{canon}_{k}$. From here, the commutant of $\kappa|\mathrm{M}_{k}$ is identified with $\mathfrak{B}(H)\otimes I_{k}$ [Con00, Section 50]. Since $\kappa|\mathfrak{A}$ takes values in this commutant, it in turn must have the form $\pi\otimes I_{k}\cong\pi^{(k)}$ for some representation $\pi$ of $\mathfrak{A}$ on $H$. Finally, if a projection of $K$ commutes with both $\kappa|\mathrm{M}_{k}$ and $\kappa|\mathfrak{A}$, then it must have the form $P\otimes I_{k}$ for some projection $P\in\pi(\mathfrak{A})^{\prime}$ (see, for instance, [Con00, Proposition 50.11]). Putting these ingredients together gives parts (a–c), and part (d) is a direct calculation. ∎
Generalizing further, given two positive integers $k$ and $\ell$, we write $\mathrm{M}_{k,\ell}(\mathfrak{A})$ for the vector space of $k$$-\ell$ matrices with entries in $\mathfrak{A}$. We multiply entries of $\mathrm{M}_{k,\ell}(\mathfrak{A})$ and $\mathrm{M}_{\ell,m}(\mathfrak{A})$ according to the usual rules of matrix multiplication. Given also a representation $\pi$ of $\mathfrak{A}$ on $H$ and an element $a=[a_{ij}]$ of $\mathrm{M}_{\ell,k}(\mathfrak{A})$, we write $\pi^{(\ell,k)}(a)$ for $[\pi(a_{ij})]_{ij}$, and interpret it as a linear operator from $H^{\oplus k}$ to $H^{\oplus\ell}$ by following the rules of matrix-vector multiplication.
@@ 4.3 Functionals on algebras of matrices
Let $\mathfrak{A}$ be a separable C*-algebra, $\pi$ be a separable representation of it, and $k$ be a positive integer. Consider a $k$-tuple $v_{1}$, …, $v_{k}$ in $H_{\pi}$, and form the single vector
@@
We collect these calculations into the following.
@@
By Lemma 4.3, any representation of $\mathrm{M}_{k}(\mathfrak{A})$ is of the form $\pi^{(k)}$ up to equivalence, and so any positive functional on $\mathrm{M}_{k}(\mathfrak{A})$ is associated to some such representation by a cyclic vector that can be written in the form in Lemma 4.4. Therefore the map in (4.4) restricts to a bijection between the closed cones $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\mathrm{M}_{k}(\mathfrak{A})_{+}^{\ast}$. With this isomorphism understood, we henceforth regard the weak topology as defined on either space.
The weak topology on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ is generally not metrizable, but, since $\mathfrak{A}$ is separable, it becomes metrizable upon restriction to any norm-bounded subset of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$. In addition, any element $\varphi$ of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ satisfies
$\|\langle\varphi,\cdot\rangle\|_{\mathrm{M}_{k}(\mathfrak{A})^{\ast}}=\langle\varphi,1\otimes I_{k}\rangle=\overline{\mathrm{tr}}_{k}\varphi(1)$ (4.5)
(see [Con00, Corollary 7.3]). This implies that the norm of $\mathrm{M}_{k}(\mathfrak{A})^{\ast}$ becomes weak continuous when we restrict it to $\mathrm{M}_{k}(\mathfrak{A})_{+}^{\ast}$ and identify that cone with $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. Therefore the restriction of the weak topology to either cone is still metrizable.
Pairings like (4.3) are a convenient way to pass between direct sums with different numbers of summands.
@@
Fix positive integers $k$ and $\ell$. Let $\pi$ be a representation of $\mathfrak{A}$, let $v_{1},\dots,v_{k}$ be a $k$-tuple in $H_{\pi}$, and let $\varphi:=\Phi^{\pi}_{v_{1},\dots,v_{k}}$. Let $a\in\mathrm{M}_{\ell,k}(\mathfrak{A})$, define
$\left[\begin{array}[]{c}y_{1}\\ \vdots\\ y_{\ell}\end{array}\right]=\pi^{(\ell,k)}(a)\left[\begin{array}[]{c}v_{1}\\ \vdots\\ v_{k}\end{array}\right]\ \in H_{\pi}^{\oplus\ell}$
and let $\psi:=\Phi^{\pi}_{y_{1},\dots,y_{\ell}}$. Then these completely positive maps are related by
$\ell\langle\psi,b\rangle:=k\langle\varphi,a^{\ast}ba\rangle\qquad(b\in\mathrm{M}_{\ell}(\mathfrak{A})).$ (4.6)
@@
This is just a matter of matrix multiplication and taking adjoints:
$\ell\langle\psi,b\rangle=\left\langle\pi^{(\ell,k)}(ba)\left[\begin{array}[]{c}v_{1}\\ \vdots\\ v_{k}\end{array}\right],\pi^{(\ell,k)}(a)\left[\begin{array}[]{c}v_{1}\\ \vdots\\ v_{k}\end{array}\right]\right\rangle=k\langle\varphi,a^{\ast}ba\rangle.$
@@
First, we extend the definitions of $\pi$-normal and $\pi$-singular to completely positive maps by simply referring to minimal dilations rather than GNS representations. Then, given a representation $\pi$ of $\mathfrak{A}$ and a map $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, we can apply Proposition 3.3 to the map $\langle\varphi,\cdot\rangle$ and representation $\pi^{(k)}$. Inspecting the resulting decomposition of $\langle\varphi,\cdot\rangle$ in the light of the isomorphism (4.4) and Lemmas 4.3 and 4.4, we obtain the following.
@@
where the right-hand terms refer to the Lebesgue decomposition of $\langle\varphi,\cdot\rangle$ with respect to $\pi$.
In particular, $\varphi$ is $\pi$-normal (resp. $\pi$-singular) if and only if $\langle\varphi,\cdot\rangle$ is $\pi^{(k)}$-normal (resp. $\pi^{(k)}$-singular). ∎
Now we consider the generalization of Proposition 3.5. Suppose that $\tau$ is a tracial positive functional on $\mathfrak{A}$, and let it be associated to its GNS representation $\lambda$ on $H$ by the cyclic vector $\xi$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and assume that it is associated to $\lambda^{\oplus\infty}$. Then Proposition 3.5 tells us that $\langle\varphi,\cdot\rangle$ is associated to $(\lambda^{(k)})^{\oplus\infty}$. However, at this point we must be careful. The representation $\lambda^{(k)}$ has a cyclic $k$-tuple given by
@@
$\lambda\otimes\mathrm{mult}_{k}\cong(\lambda^{(k)})^{\oplus k}$
of the algebra $\mathrm{M}_{k}(\mathfrak{A})=\mathfrak{A}\otimes\mathrm{M}_{k}$ on $H\otimes\mathrm{M}_{k}$. This has a cyclic vector given by
$\Xi=\xi\otimes I_{k}=[\xi_{1},\dots,\xi_{k}],$
@@
We have reached a source of possible confusion. Suppose that $\varphi$ is associated to $\lambda^{\oplus k}$ by the vectors $x_{i}=T\xi_{i}$, $i=1,\dots,k$, as in Proposition 4.7. Suppose further that we define another completely positive map $\psi:\mathfrak{A}\to\mathrm{M}_{k}$ by the pairing relationship
$\langle\psi,b\rangle=\langle\varphi,a^{\ast}ba\rangle\qquad(b\in\mathrm{M}_{k}(\mathfrak{A}))$
for some $a=[a_{ij}]\in\mathrm{M}_{k}(\mathfrak{A})$, as in Lemma 4.5. Then $\psi$ is associated to the tuple of vectors $y_{1}$, …, $y_{k}\in H^{\oplus k}$ given by the formula
$\left[\begin{array}[]{c}y_{1}\\ \vdots\\ y_{k}\end{array}\right]=[\lambda^{\oplus k}(a_{ij})]\left[\begin{array}[]{c}x_{1}\\ \vdots\\ x_{k}\end{array}\right].$
Both sides of this formula are column vectors of height $k$ with individual entries that are elements of $H^{\oplus k}$ – that is, they are $k$-column vectors of $k$-column vectors. Correspondingly, $[\lambda^{\oplus k}(a_{ij})]$ is a $k$-by-$k$ matrix of operators with individual entries that are $k$-fold inflations.
So there are two $k$-fold inflations at work here, and it is important to keep their roles separate. The first inflation is needed in general to prove enough ‘room’ in the representation $\lambda^{\oplus k}$ to find a whole $k$-tuple $x_{1}$, …, $x_{k}$ that associates $\varphi$ to $\lambda^{\oplus k}$. The second inflation is needed to express how a $k$-by-$k$ matrix of elements of $\mathfrak{A}$ acts to turn $\varphi$ into $\psi$.
@@
With Proposition 4.7 in hand, we finally reach the definition that makes full sense of Theorems A and C. To explain it, recall that $\lambda^{\oplus k}(\mathfrak{A})^{\prime}$ has a natural identification with $\mathrm{M}_{k}(\lambda(\mathfrak{A})^{\prime})$, which allows us to regard $\widetilde{\tau}\otimes\mathrm{tr}_{k}$ as a normal tracial positive functional on either algebra.
@@
Let $\lambda$ and $\tau$ be as above. If $\varphi$ is $\lambda$-normal, the its Fuglede–Kadison determinant with respect to $\tau$ is

### v3

@@ 4.1 Tuples of vectors and completely positive maps
Let $\mathfrak{A}$ be a separable C*-algebra, $\pi$ be a separable representation of it, and $k$ be a positive integer. Let $v_{1}$, …, $v_{k}\in H_{\pi}$, and regard the tuple $V:=[v_{1},\dots,v_{k}]$ as a linear map from $\mathbf{C}^{\oplus k}$ to $H_{\pi}$. To keep track of how these vectors move together under the action of $\pi$, we can consider the $\mathrm{M}_{k}$-valued function
@@
If $\varphi$ is a continuous linear map from $\mathfrak{A}$ to $\mathrm{M}_{k}$, then a classic theorem of Stinespring asserts that $\varphi$ equals $\Phi^{\pi}_{V}$ for some representation $\pi$ and $k$-tuple $V$ if and only if $\varphi$ is completely positive. See, for example, [139] or [38, Theorem 34.7]. The most substantial part of this result is the construction of $\pi$ and $V$ from an abstract completely positive map $\varphi$. This is an $\mathrm{M}_{k}$-valued generalization of the GNS construction. If we require in addition that $V$ be cyclic for $\pi$, then the resulting pair $(\pi,V)$ is unique up to unitary equivalence. Its underlying representation is called the minimal dilation of $\varphi$ and is denoted by $\pi_{\varphi}$. Many assertions about GNS representations adapt straightforwardly to minimal dilations, and we extend our terminology accordingly. For example, two matrix-valued completely positive maps are disjoint if their minimal dilations are disjoint.
Generalizing our usage in the scalar-valued case, we write that the completely positive map in (4.1) is associated to $\pi$ by $V$, or we call $\Phi^{\pi}_{V}$ the type of $V$ in $\pi$.
In the present paper we do not use the property of complete positivity directly, but we adopt this as the standard term for the class of maps we need to work with. We write $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ for the space of all continuous linear maps from $\mathfrak{A}$ to $\mathrm{M}_{k}$, and $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ for the subset of all completely positive ones. Complete positivity is defined by a family of closed linear inequalities, so $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ is a weak-closed cone in $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$.
A completely positive map $\varphi:\mathfrak{A}\to\mathrm{M}_{k}$ is normalized if $\overline{\mathrm{tr}}_{k}\varphi(1)=1$, and unital if $\varphi(1)=I_{k}$. A tuple of vectors in $H_{\pi}$ is orthonormal if and only if their type is unital.
@@
The vector space $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ has a natural topology obtained by applying the weak topology in each matrix entry. Henceforth we simply refer to this as ‘the weak topology’ of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$. As a vector space topology, it also defines a uniform structure on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ (see, for instance, [54, Section 8.1] for the basics of uniform structures). We need this uniform structure a few times through the next lemma.
@@
These properties are elementary for the inner product map $H_{\pi}\times H_{\pi}\to\mathbf{C}$. They follow for types by arguing pointwise for each $i$, $j\in\{1,\dots,k\}$, and ${a\in\mathfrak{A}}$. ∎
Consider again a representation $\pi$ and a tuple $v_{1}$, …, $v_{k}$ in $H_{\pi}$. If $a=[a_{ij}]$ is an $\ell$-by-$k$ matrix of elements of $\mathfrak{A}$, then we can define a new $\ell$-tuple in $H_{\pi}$ by the formula
$\left[\begin{array}[]{c}y_{1}\\ \vdots\\ y_{\ell}\end{array}\right]:=[\pi(a_{ij})]\cdot\left[\begin{array}[]{c}v_{1}\\ \vdots\\ v_{k}\end{array}\right],$ (4.2)
understood by following the rules of matrix-vector multiplication. For example, if $a=[q_{ij}\cdot 1]$ for some scalar matrix $Q=[q_{ij}]$, then we can identify $[\pi(a_{ij})]$ with $I_{H_{\pi}}\otimes Q$, and (4.2) becomes
$[y_{1},\dots,y_{\ell}]^{\mathrm{T}}:=(I_{H_{\pi}}\otimes Q)[v_{1},\dots,v_{k}]^{\mathrm{T}}.$ (4.3)
If the tuples $v_{1}$, …, $v_{k}$ and $y_{1}$, …, $y_{\ell}$ are related as in (4.2), and $\varphi$ and $\psi$ are their respective types, then $\psi$ may be written using the tuple $v_{1}$, …, $v_{k}$ like this:
[…]
Writing this right-hand side in terms of $\varphi$ itself yields the following.
Lemma 4.3.
In the situation above, we have
[…]
As a result, with $[a_{ij}]$ held fixed, $\psi$ is continuous as a function of $\varphi$ for the weak topologies.
In particular, in the special case of (4.3), we have
$\psi(b):=(Q^{\mathrm{T}})^{\ast}\varphi(b)Q^{\mathrm{T}}\qquad(b\in\mathfrak{A}).$ (4.5)
Sometimes we need to consider all the ways in which two matrix-valued completely positive maps could ‘sit together’ inside a larger one. To describe these, we borrow a term from Furstenberg’s classic work [65] in ergodic theory. Let $k$ and $\ell$ be positive integers, and let
$K\>=\{1,\dots,k\}\qquad\hbox{and}\qquad L:=\{k+1,\dots,k+\ell\}.$
Definition 4.4.
In particular, the diagonal joining is defined by
This terminology is not standard in representation theory, but it is a convenient way to organize various arguments below.
If $V=[v_{1},\dots,v_{k}]$ and $W=[w_{1},\dots,w_{\ell}]$ are two tuples in a representation $\pi$, then the combined type $\Phi^{\pi}_{[V,W]}$ is a joining of $\Phi^{\pi}_{V}$ and $\Phi^{\pi}_{W}$. These two tuples generate orthogonal subrepresentations of $\pi$ if and only if
$\Phi^{\pi}_{[V,W]}=\mathrm{diag}(\Phi^{\pi}_{V},\Phi^{\pi}_{W}).$
On the other hand, given any joining $\theta$ of $\Phi^{\pi}_{V}$ and $\Phi^{\pi}_{W}$, the minimal dilation $\pi_{\theta}$ contains canonical copies of both tuples. We may therefore characterize disjointness as follows.
Lemma 4.5.
If $\varphi:\mathfrak{A}\to\mathrm{M}_{k}$ and $\psi:\mathfrak{A}\to\mathrm{M}_{\ell}$ are completely positive, then they are disjoint if and only if $\mathrm{diag}(\varphi,\psi)$ is their only joining. ∎
In fact, starting with Furstenberg’s paper [65], the uniqueness of the product joining is taken as the definition of ‘disjointness’ in ergodic theory, where the lack of orthogonal complements makes other definitions impractical or senseless.
@@ 4.2 Matrices over algebras
Most of the results about positive functionals in Section 3 have standard generalizations to $\mathrm{M}_{k}$-valued completely positive maps, and the proofs generally run along the same lines. Alternatively, many facts about $\mathrm{M}_{k}$-valued completely positive maps can be reduced directly to the scalar-valued case by regarding tuples of vectors in a representation of $\mathfrak{A}$ as single vectors in a representation of an algebra of matrices over $\mathfrak{A}$.
If $\mathfrak{A}$ is a C-algebra and $k$ is a positive integer, then we write $\mathrm{M}_{k}(\mathfrak{A})$ for the algebra of $k$-by-$k$ matrices with entries from $\mathfrak{A}$. The algebra operations combine those of $\mathfrak{A}$ with the usual rules for matrices, and we define an involution on $\mathrm{M}_{k}(\mathfrak{A})$ by transposing and applying the involution of $\mathfrak{A}$ entry-wise. Then $\mathrm{M}_{k}(\mathfrak{A})$ is a separable C*-algebra in a canonical way. It has a natural identification with $\mathfrak{A}\otimes\mathrm{M}_{k}$, and we pass freely between these two points of view. See, for instance, [38, Section 34].
If $\pi$ is a representation of $\mathfrak{A}$ and $k$ is a positive integer, then we define a representation $\pi^{(k)}$ of $\mathrm{M}_{k}(\mathfrak{A})$ on $H_{\pi}^{\oplus k}$ by setting $\pi^{(k)}([a_{ij}]):=[\pi(a_{ij})]$ and following the rules of matrix-vector multiplication as in (4.2). From another point of view, we can identify $H_{\pi}^{\oplus k}$ with $H_{\pi}\otimes\mathbf{C}^{\oplus k}$, and then $\pi^{(k)}$ is the Kronecker product of $\pi$ with the canonical representation
@@
    We have $\pi^{(k)}\otimes I_{\ell}\simeq(\pi\otimes I_{\ell})^{(k)}$ for every positive integer $k$ and {${}\ell\in\{1,2,\dots,\}\cup\{\infty\}$.
@@
Any representation $\kappa$ of $\mathrm{M}_{k}(\mathfrak{A})\cong\mathfrak{A}\otimes\mathrm{M}_{k}$ on a Hilbert space $K$ is generated by the commuting representations $\kappa|\mathrm{M}_{k}$ and $\kappa|\mathfrak{A}$. For $\mathrm{M}_{k}$, every representation is an inflation of $\mathrm{canon}_{k}$: this is a classical result of pure algebra, or a special case of the representation theory of C*-algebras of compact operators [38, Section 16]. We can therefore write $K$ as $H^{\oplus k}$ for some auxiliary Hilbert space $H$ so that $\kappa|\mathrm{M}_{k}$ is identified with $I_{H}\otimes\mathrm{canon}_{k}$. From here, the commutant of $\kappa|\mathrm{M}_{k}$ is identified with $\mathfrak{B}(H)\otimes I_{k}$ [38, Section 50]. Since $\kappa|\mathfrak{A}$ takes values in this commutant, it in turn must have the form $\pi\otimes I_{k}\simeq\pi^{(k)}$ for some representation $\pi$ of $\mathfrak{A}$ on $H$. Finally, if a projection of $K$ commutes with both $\kappa|\mathrm{M}_{k}$ and $\kappa|\mathfrak{A}$, then it must have the form $P\otimes I_{k}$ for some projection $P\in\pi(\mathfrak{A})^{\prime}$ [38, Proposition 50.11]. Putting these ingredients together gives parts (a–c), and part (d) is a direct calculation. ∎
@@ 4.3 Functionals on algebras of matrices
Let $\pi$ be a representation of $\mathfrak{A}$ as before. Consider a $k$-tuple $v_{1}$, …, $v_{k}$ in $H_{\pi}$, and form the single vector
@@
These calculations have the following immediate consequences.
@@
With this isomorphism understood, we henceforth regard either space as carrying the dual norm or weak topology.
By Lemma 4.6, any representation of $\mathrm{M}_{k}(\mathfrak{A})$ is of the form $\pi^{(k)}$ up to equivalence. So any positive functional on $\mathrm{M}_{k}(\mathfrak{A})$ is associated to such a representation by a cyclic vector that can be written in the form in Lemma 4.7. It follows that the map in (4.8) restricts to a bijection between the closed cones $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\mathrm{M}_{k}(\mathfrak{A})_{+}^{\ast}$. Because of positivity, the restriction of the dual norm to these cones satisfies
$\|\langle\varphi,\cdot\rangle\|_{\mathrm{M}_{k}(\mathfrak{A})^{\ast}}=\langle\varphi,1\otimes I_{k}\rangle=\overline{\mathrm{tr}}_{k}\varphi(1)\qquad(\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+})$ (4.9)
(see [38, Corollary 7.3]).
The weak topology on $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ is generally not metrizable, or even first countable. However, because $\mathfrak{A}$ is separable, this topology becomes metrizable upon restriction to any subset of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$ that is bounded in the dual norm [37, Theorem V.5.1]. The next lemma lies between these cases. It must surely be classical, but I have not found a reference for it.
@@
The restriction of the weak topology to $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ is locally compact and second countable.
@@
By considering pairing functionals on $\mathrm{M}_{k}(\mathfrak{A})$ as in (4.7) instead of positive definite maps on $\mathfrak{A}$, we may reduce to the case $k=1$. Having done so, let
$U_{r}:=\{\varphi\in\mathfrak{A}^{\ast}_{+}:\ \varphi(1)<r\}\qquad(r>0).$
Each of these sets is relatively weak-open in $\mathfrak{A}^{\ast}_{+}$. On the other hand, by (4.9), the weak-closure $\overline{U_{r}}$ is equal to the intersection of $\mathfrak{A}^{\ast}_{+}$ with a closed ball of radius $r$ in $\mathfrak{A}^{\ast}$. Therefore $\overline{U_{r}}$ is metrizable and also compact, the latter by the Banach–Alaoglu theorem, and hence it is second countable. So $\mathfrak{A}^{\ast}_{+}$ is covered by the sequence $U_{1}$, $U_{2}$, …of open subsets, each of which is precompact and second countable in the weak topology. ∎
@@
First, we extend the definitions of $\pi$-normal and $\pi$-singular to completely positive maps by simply referring to minimal dilations rather than GNS representations. Then, given a representation $\pi$ of $\mathfrak{A}$ and a map $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, we can apply Proposition 3.3 to the map $\langle\varphi,\cdot\rangle$ and representation $\pi^{(k)}$. Inspecting the resulting decomposition of $\langle\varphi,\cdot\rangle$ in the light of the isomorphism (4.8) and Lemmas 4.6 and 4.7, we obtain the following.
@@
where the right-hand terms refer to the Lebesgue decomposition of $\langle\varphi,\cdot\rangle$ with respect to $\pi$. The summands are linear in $\varphi$.
In particular, $\varphi$ is $\pi$-normal (respectively, $\pi$-singular) if and only if $\langle\varphi,\cdot\rangle$ is $\pi^{(k)}$-normal (respectively, $\pi^{(k)}$-singular). ∎
Now we consider the generalization of Proposition 3.5. Suppose that $\tau$ is a tracial positive functional on $\mathfrak{A}$, and let it be associated to its GNS representation $\lambda$ by the cyclic vector $\xi$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and assume that it is associated to $\lambda^{\oplus\infty}$. By Lemma 4.7 and Lemma 4.6(d), $\langle\varphi,\cdot\rangle$ is associated to $(\lambda^{(k)})^{\oplus\infty}$. However, at this point we must be careful. The representation $\lambda^{(k)}$ has a cyclic $k$-tuple given by
@@
$\lambda\otimes\mathrm{mult}_{k}\simeq(\lambda^{(k)})^{\oplus k}$
of the algebra $\mathrm{M}_{k}(\mathfrak{A})=\mathfrak{A}\otimes\mathrm{M}_{k}$ on $H_{\lambda}\otimes\mathrm{M}_{k}$. This has the single cyclic vector
$\Xi:=\xi\otimes I_{k}=[\xi_{1},\dots,\xi_{k}],$
@@
We have reached a source of possible confusion. Suppose that $\varphi$ is associated to $\lambda^{\oplus k}$ by the vectors $x_{i}=T\xi_{i}$, $i=1,\dots,k$, as in Proposition 4.10. Suppose further that $a=[a_{ij}]\in\mathrm{M}_{k}(\mathfrak{A})$ and that we define a new tuple $y_{1}$, …, $y_{k}$ by (4.2), so that the type $\psi$ of this new tuple is given by (4.4). In this case, both sides of (4.2) are column vectors of height $k$ with individual entries that are elements of $H_{\lambda}^{\oplus k}$ – that is, they are column $k$-vectors of column $k$-vectors. Correspondingly, $[\lambda^{\oplus k}(a_{ij})]$ is a $k$-by-$k$ matrix of operators with individual entries that are $k$-fold inflations.
So there are two $k$-fold inflations at work here, and it is important to keep their roles separate. The first inflation provides enough ‘room’ in the representation $\lambda^{\oplus k}$ to find a whole $k$-tuple $x_{1}$, …, $x_{k}$ that associates $\varphi$ to $\lambda^{\oplus k}$. The second inflation is needed to express how a $k$-by-$k$ matrix of elements of $\mathfrak{A}$ acts to turn $\varphi$ into $\psi$.
@@
With Proposition 4.10 in hand, we finally reach the definition that makes full sense of Theorems AB, and C. To explain it, recall that $\lambda^{\oplus k}(\mathfrak{A})^{\prime}$ has a natural identification with $\mathrm{M}_{k}(\lambda(\mathfrak{A})^{\prime})$, which allows us to regard $\widetilde{\tau}\otimes\mathrm{tr}_{k}$ as a normal tracial positive functional on either algebra.
@@
Let $\lambda$ and $\tau$ be as above. If $\varphi$ is $\lambda$-normal, then its Fuglede–Kadison determinant with respect to $\tau$ is